What are the dimensions in SI units of a square-shaped acre of ground if 1.000 square mile = 6.40 times 10^2 acres? Note that 1.00 km = 0.621 mi. Ensure that your answer has the correct number of significant figures.

____________________m
Answer #1

1.00 km = 0.621 mi

1 mi = (1.00/0.621) km

1 square mile = 6.40 x 10^2 acres

1 acre = \frac{1}{6.40*10^2} square miles

for a square shaped acre, dimension of one side

= \frac{1}{\sqrt{6.40*10^2}}\ miles

= \frac{1}{\sqrt{6.40*10^2}}*\frac{1.00}{0.621}*1000\ m

=63.653\ m = 63.7\ m         (least no of significant digits is 3)

Dimensions of ground are 63.7 m x 63.7 m